# {tocentry} Vibrational Energy Dissipation in Non-Contact Single-Molecule Junctions Governed by Local Geometry and Electronic Structure

### v1

@@ Abstract
The vibrational dynamics of adsorbate molecules in single-molecule junctions depend critically on the geometric structure and electronic interactions between molecule and substrate. Vibrations, excited mechanochemically or by external stimuli, dissipate energy into substrate electrons and phonons. Energy dissipation leads to the broadening of spectral lines, vibrational lifetimes, and the coupling between molecular and substrate phonons. It affects molecular manipulation, giving rise to nanoscale friction, and contributes to scanning probe and surface spectroscopy signals. We present an approach to disentangle adsorbate vibrational dynamics in non-contact junctions by employing density functional theory, machine learning, and non-adiabatic molecular dynamics. Focusing on the CO-functionalised Cu surfaces representing a single-molecule junction, a widely studied system in scanning probe and energy dissipation experiments, we reveal strong vibrational mode specificity governed by the interplay of electron-phonon and phonon-phonon coupling. Electron-phonon relaxation rates vary by two orders of magnitude between modes and sensitively depend on the tip-substrate geometry. We find evidence of a weak non-additive effect between both energy dissipation channels, where electron-phonon coupling enhances phonon-phonon coupling. Our predicted vibrational lifetimes agree with infrared spectroscopy and helium scattering experiments. Finally, we outline how our findings can inform and enhance scanning probe experiments.
keywords:
mechanochemistry, single-molecule microscopy, electron-phonon coupling, anharmonic phonon coupling, density functional theory, vibrational energy dissipation1 Introduction
The vibrational dynamics of atoms and molecules adsorbed on a surface provide valuable insights into the interactions and energy exchange between adsorbate and substrate. These dynamics are governed by interactions between adsorbates and the surface and quantum-mechanical energy dissipation mechanisms,1, 2 Energy dissipation mechanisms in non-contact single-molecule junctions significantly affect measured signals in atomic force microscopy (AFM), lateral force microscopy (LFM),3, 4 scanning tunnelling microscopy (STM),5 inelastic electron tunnelling spectroscopy (IETS), and infrared spectroscopy experiments.6, 7 In such experiments, a key challenge is to disentangle geometric structure, dynamics, and energy dissipation mechanisms in the measured signal: For example, AFM-induced hopping of atoms and molecules leads to measurable energy dissipation,8, 9, 10, 3 yet the underlying dissipation pathways remain unclear. Helium atom scattering (HAS) experiments of CO adsorbed on Cu surfaces have found a strong dependence of vibration frequencies on both the surface facet and the presence of defects.11
@@
Figure 1 shows a schematic setup of an idealised single-molecule junction as it would be present in an AFM experiment. If vibrations of the molecule in the junction are excited, for example, through the motion of the tip, energy dissipation away from the molecule can occur by excitation of phonons or electron-hole-pair (EHP) excitations in the metal tip or substrate. The former is due to anharmonicity that induces phonon-phonon coupling between adsorbate and substrate.2 The latter is due to coupling between electronic and nuclear motions, commonly referred to as electron-phonon coupling.6 Intramolecular vibrational relaxation and radiative processes, such as infrared emission, also contribute to energy dissipation, though their role is typically minor for small adsorbates.15, 16 However, for larger adsorbates, intramolecular vibrational relaxation can become a significant dissipation channel.17 Consequently, it is of great interest to explore how the geometric structure of a surface or junction affects the mechanical and electronic interactions with the adsorbate, which govern the relative contributions of electron-phonon and phonon-phonon coupling to the vibrational dynamics. Gaining control over these energy dissipation channels requires the ability to identify their distinct signatures and disentangle their interplay—an essential step towards rational design of surface functionality at the atomic scale.
@@
In this work, we study competing energy dissipation mechanisms and their effect on the vibrational dynamics of adsorbate molecules for the exemplary case of CO-decorated tip-surface copper junctions. This system has been extensively characterised by experiment18, 6, 19 and is widely used in atomic-scale resolution scanning probe measurements.20 Various previous theoretical studies of energy dissipation for CO on different, idealised Cu surfaces21, 22, 23 have been reported, but no clear analysis of the role of the local adsorption geometry exists to date. A specific challenge in this system for electronic structure theory is the correct description of the potential energy surface of CO on Cu(111).24, 25, 26 We address this by using the state-of-the-art screened hybrid functional HSE0627 in combination with the MBD-NL van der Waals (vdW) correction28 to correctly describe the adsorption site and the potential energy surface. From a first-principles perspective, we describe the details and main energy dissipation mechanisms, namely electron-phonon and phonon-phonon coupling, which govern the dynamics at the atomic scale.2 We find that the surface geometry plays a significant role in these mechanisms, with flatter surfaces exhibiting stronger electron-phonon and phonon-phonon coupling. Additionally, we observe a strong mode dependence of electron-phonon coupling, with relaxation rates differing by two orders of magnitude. The molecular dynamics with electronic friction approach29, 21 allows us to study the interplay of the two dissipation channels and their concerted effects on the relaxation dynamics. We find that the two effects display weak non-additivity, suggesting that subtle dynamical steering may arise from one or both mechanisms. Our results are in good agreement with existing experiments and support the interpretation of single-molecule dynamics in non-contact junctions.
@@ 2 Results
2.1 Models of tip geometry in non-contact junctions
As shown in Figure 1, we can consider two interaction regimes for a non-contact CO tip-molecule-surface junction: In one case, the molecule is attached to the surface, in the other case, it is attached to the tip. The molecule can be excited by a variety of processes: The motion of the tip will displace the molecule with respect to the corrugated potential energy landscape of the underlying substrate.30 The mechanochemical forces the molecule experiences in the process will excite its vibrations. The molecular vibrations can absorb light,6 supported by field enhancement in the junction.31 A potential bias and electron tunnelling can lead to current-induced forces.3 In all cases, molecular vibrations will be excited, leading to energy dissipation into EHPs and phonons in the metal. The case where the molecule is attached to the surface is not too different from the scenario in the absence of the tip. Due to multiple Angstroms of distance between the tip and the molecule, mechanochemical forces on the molecule will be weak. Other perturbation scenarios have previously been studied.23, 22 For simplicity, we consider here only surfaces in both the slab and slab–adatom systems that correspond to the Cu(111) facet.
We turn our attention to the second case, where the molecule is attached to the tip. Regardless of the origin of the vibrational excitation, energy dissipation into the underlying substrate in this scenario will be almost negligible, as evidenced by calculations for CO on Cu in Section S1 of the Supporting Information. Therefore, molecular vibrations will dominantly dissipate energy into EHPs and phonons in the tip. This motivates our model design: We consider four different possible “tip” geometries: CO on a sharp tip (sharp tip), CO on a blunt tip (blunt tip), CO on an adatom on a surface slab (slab adatom), and CO on a slab (slab), which are depicted in Figure 1c and Figure S3 in the Supporting Information. The underlying surface, in the case of lateral motion of the tip, only serves as a corrugated energy landscape that triggers molecular vibrations4 and therefore does not have to be explicitly considered in the atomistic model. The “tip” models are therefore represented by two-dimensional periodic surface slab geometries. Further details on the models and the computational methodology are provided in the Methods section and Section S1 of the Supporting Information.
@@ 2.2 Vibrational mode analysis
CO adsorbed on a flat metal surface has four distinct non-degenerate vibrational modes: frustrated translation (FT), frustrated rotation (FR), surface-adsorbate stretch (SA), and the internal stretch of the CO bond (IS), as illustrated in Figure 2. The FT and FR modes consist of two degenerate modes each.
We determine vibrational energies from second-order force constants (i.e. the Hessian) for all atoms in the molecule and the top layers of the four surface models. Simulations include ten layers of freely movable Cu atoms, spanning hundreds of degrees of freedom. Energies, forces, and second-order force constants are evaluated with a MACE graph neural network model32 trained on first-principles data (see Methods). The MACE potential accurately reproduces the underlying HSE06 data, correctly predicting the adsorption sites: the top site for CO on a (111) surface slab of copper and the hollow site for the adatom in the CO-adatom arrangement. (see Section S2 of the Supporting Information). The calculated vibrational frequencies for CO on the Cu(111) slab show good agreement with experimental data, as presented in Table 1. Interestingly, the stiffness of low-energy modes increases as the tip becomes blunter, as indicated in Table 1. Moreover, there is a jump in stiffness when transitioning from CO adsorbed on an adatom to CO adsorbed directly on the surface. HAS experiments estimated the energy of the FT modes on a flat Cu(111) surface to be , while the energy dropped to at step edges of the Cu(211) and Cu(511) surfaces.11 This trend is in good agreement with our finding that more tapered (or corrugated) surface geometries exhibit lower vibrational energies for the FT mode.
@@
The vibrational density of states (VDOS) (see Figure 3) provides an overview of all vibrational modes. We differentiate between modes localised dominantly at the molecule (orange in Figure 3), where the majority of the kinetic energy is contained in the CO and substrate modes (black in Figure 3), where most of the energy is in the motion of the Cu atoms. The FR, SA, and IS vibrational modes of CO exhibit higher frequencies compared to most copper substrate modes, which typically range between approximately and . We note that the VDOS shown here is normalised by the number of atoms. The substrate vibrational states are more localised in energy (or frequency) for slab and slab-adatom surfaces and more delocalised for the sharp and blunt tips. For sharp and blunt tips, a significant portion of the surrounding space is vacuum, whereas this space is occupied by atoms in the slab and slab-adatom surfaces. The presence of more atoms in these flatter configurations results in a greater number of substrate vibrational states per volume. This facilitates stronger overall coupling between the CO and substrate vibrational modes, provided that the CO and Cu degrees of freedom are sufficiently coupled. As a rule of thumb, if the off-diagonal elements of the Hessian (for CO-Cu coupling) are of a similar magnitude, we can expect strong coupling. As shown in Figure S4 in the Supporting Information, this is indeed the case. In the case of the sharp and blunt tip, we find a strong coupling for the CO to a small number of Cu degrees of freedom. The slab adatom and slab geometries exhibit weaker coupling but a larger number of states. When discussing phonon-phonon coupling, we will return to this observation.
@@ 2.3 Phonon-phonon coupling
The atomic motion of the CO molecule is coupled to the atomic motion of the Cu tip through phonon-phonon interactions and to EHP through electron-phonon coupling.2 Even at low temperatures, phonon-phonon coupling plays a significant role in the dynamics of molecular vibrations. In our study, we use a temperature of , a condition commonly employed in experiments such as AFM measurements with CO-decorated tips. Phonon-phonon coupling is determined by analysing the dynamics of a CO molecule on surface geometries where approximately 500 Cu atoms are free to move per unit cell, allowing energy transfer between CO modes and substrate modes. To determine coupling strengths, we use vibrational analysis of molecular dynamics trajectories in thermodynamic equilibrium.33 We first thermalise each system at a temperature of and then allow it to evolve in thermodynamic equilibrium over time. Over time, the velocity time series of the atoms becomes decorrelated, as the atoms transfer mechanical energy between each other. The rate of this decorrelation can be found in the cross-correlation function. For this analysis, we employ normal mode decomposition to project the velocity time series onto the basis defined by the vibrational eigenvectors. For the velocity time series of two modes and the velocity cross-correlation function is as follows:
@@
The cross-spectrum provides valuable insights into vibrational lifetimes and relative coupling strengths. Let us first discuss the coupling strength, which can be determined by analysing the off-diagonal elements . These exhibit two Lorentzian-shaped peaks at the characteristic frequencies of modes and , with their heights reflecting the coupling strength between these modes (see Section S5 of the Supporting Information). To further understand this interaction, we examine the coupling density , defined as the sum of the cross spectra for all modes correlated with the mode of interest :
A broad peak in the coupling density at a given frequency means that modes at this frequency are correlated with the mode of interest, , indicating strong vibrational coupling. Our analysis reveals that the coupling is strongest with adjacent substrate vibrational modes, as shown in Figure 4 and Figures S18, S19, S20, and S21 in the Supporting Information. This effect is less pronounced in case of energetically decoupled SA and IS modes. The coupling between different CO modes is small in comparison, indicating that intramolecular vibrational redistribution plays only a minor role. Notably, the FR and IS modes exhibit the strongest coupling.
The geometry of the surface has a substantial impact on the coupling density. Among the different geometries, the sharp tip exhibits the weakest coupling, while the slab geometry demonstrates the strongest coupling density. This observation can be attributed to the VDOS discussed earlier. As surface geometry becomes more atomically flat, the VDOS increases, offering more vibrational states to accommodate the energy of CO modes and thereby enhancing phonon-phonon coupling.

### v2

@@ Abstract
The vibrational dynamics of adsorbate molecules in single-molecule junctions depend critically on the geometric structure and electronic interactions between molecule and substrate. Vibrations, excited mechanochemically or by external stimuli, dissipate energy into substrate electrons and phonons. Energy dissipation leads to the broadening of spectral lines, vibrational lifetimes, and the coupling between molecular and substrate phonons. It affects molecular manipulation, giving rise to nanoscale friction, and contributes to scanning probe and surface spectroscopy signals. We present an approach to disentangle adsorbate vibrational dynamics in non-contact junctions by employing Density Functional Theory, machine learning, and non-adiabatic molecular dynamics. Focusing on the CO-functionalised Cu surfaces representing a single-molecule junction, a widely studied system in scanning probe and energy dissipation experiments, we reveal strong vibrational mode specificity governed by the interplay of electron-phonon and phonon-phonon coupling. Electron-phonon relaxation rates vary by two orders of magnitude between modes and sensitively depend on the tip-substrate geometry. We find evidence of a weak non-additive effect between both energy dissipation channels, where electron-phonon coupling enhances phonon-phonon coupling. Our predicted vibrational lifetimes agree with infrared spectroscopy and helium scattering experiments. Finally, we outline how our findings can inform and enhance spectroscopy and scanning probe experiments.
Keywords: Nanoscale energy dissipation, vibrational spectroscopy, nanomechanics, electron-phonon coupling, phonon-phonon coupling, electronic structure theory
1 Introduction
The vibrational dynamics of atoms and molecules adsorbed on a surface provide valuable insights into the interactions and energy exchange between adsorbate and substrate. These dynamics are governed by interactions between adsorbates and the surface and quantum-mechanical energy dissipation mechanisms.1, 2 Energy dissipation mechanisms in non-contact single-molecule junctions significantly affect measured signals in atomic force microscopy (AFM), lateral force microscopy (LFM),3, 4 scanning tunnelling microscopy (STM),5 inelastic electron tunnelling spectroscopy (IETS), and infrared spectroscopy experiments.6, 7 In such experiments, a key challenge is to disentangle geometric structure, dynamics, and energy dissipation mechanisms in the measured signal: For example, AFM-induced hopping of atoms and molecules leads to measurable energy dissipation,8, 9, 10, 3 yet the underlying dissipation pathways remain unclear. Helium atom scattering (HAS) experiments of CO adsorbed on Cu surfaces have found a strong dependence of vibrational frequencies on both the surface facet and the presence of defects.11
@@ 1 Introduction
Figure 1 shows a schematic setup of an idealised single-molecule junction as it would be present in an AFM experiment. If vibrations of the molecule in the junction are excited, for example, through the motion of the tip, energy dissipation away from the molecule can occur by excitation of phonons or electron-hole-pair (EHP) excitations in the metal tip or substrate. The former is due to anharmonicity that induces phonon-phonon coupling (PPC) between adsorbate and substrate.2 The latter is due to coupling between electronic and nuclear motions, commonly referred to as electron-phonon coupling (EPC).6 Intramolecular vibrational relaxation and radiative processes, such as infrared emission, also contribute to energy dissipation, though their role is typically minor for small adsorbates.15, 16 However, for larger adsorbates, intramolecular vibrational relaxation can become a significant dissipation channel.17 Consequently, it is of great interest to explore how the geometric structure of a surface or junction affects the mechanical and electronic interactions with the adsorbate, which govern the relative contributions of EPC and PPC to the vibrational dynamics. Gaining control over these energy dissipation channels requires the ability to identify their distinct signatures and disentangle their interplay—an essential step towards rational design of surface functionality at the atomic scale.
@@
In this work, we study competing energy dissipation mechanisms and their effect on the vibrational dynamics of adsorbate molecules for the exemplary case of CO-decorated tip-surface copper junctions. CO on Cu is a prototypical surface-adsorbate system. It serves as a benchmark in surface science, enabling the systematic investigation of fundamental processes such as adsorption, diffusion, and vibrational energy dissipation.18, 6, 19, 20, 3 Moreover, it is widely used in atomic-resolution scanning probe measurements.21 Owing to the similar electronic structure of coinage metals, insights from CO on Cu are transferable to Ag and Au surfaces, making it relevant for experiments and applications in catalysis and energy conversion, where electron–phonon and phonon–phonon coupling govern surface reactivity.22 Various previous theoretical studies of energy dissipation for CO on different, idealised Cu surfaces23, 24, 25 have been reported, but no clear analysis of the role of the local adsorption geometry exists to date. A specific challenge in this system for electronic structure theory is the correct description of the potential energy surface of CO on Cu(111).26, 27, 28 We address this by using the state-of-the-art screened hybrid functional HSE0629 in combination with the MBD-NL van der Waals (vdW) correction30. From a first-principles perspective, we describe the details and main energy dissipation mechanisms, namely EPC and PPC, which govern the dynamics at the atomic scale.2 We find that the surface geometry plays a significant role in these mechanisms, with flatter surfaces exhibiting stronger EPC and PPC. Additionally, we observe a strong mode dependence of EPC, with relaxation rates differing by two orders of magnitude. The molecular dynamics with electronic friction (MDEF) approach31, 23 allows us to study the interplay of the two dissipation channels and their concerted effects on the relaxation dynamics. We find signs of weak non-additivity between the two effects, suggesting that subtle dynamical steering may arise from one or both mechanisms. Our results are in good agreement with existing experiments and support the interpretation of single-molecule dynamics in non-contact junctions.
@@ 2 Results
2.1 Models of the tip geometry in non-contact junctions
As shown in Figure 1, we can consider two interaction regimes for a non-contact CO tip-molecule-surface junction: In one case, the molecule is attached to the surface, in the other case, it is attached to the tip. The molecule can be excited by a variety of processes: The motion of the tip will displace the molecule with respect to the corrugated potential energy landscape of the underlying substrate.32 The mechanochemical forces the molecule experiences in the process will excite its vibrations. The molecular vibrations can absorb light,6 supported by field enhancement in the junction.33 A potential bias and electron tunnelling can lead to current-induced forces.3 In all cases, molecular vibrations will be excited, leading to energy dissipation into EHPs and phonons in the metal. The case where the molecule is attached to the surface is not too different from the scenario in the absence of the tip. The tip and molecule are separated by a distance of multiple Angstroms, so mechanochemical forces on the molecule will be weak. Other perturbation scenarios have previously been studied.25, 24 For simplicity, we consider only surfaces in both the slab and slab–adatom systems that correspond to the Cu(111) facet.
We turn our attention to the second case, where the molecule is attached to the tip. Regardless of the origin of the vibrational excitation, energy dissipation into the underlying substrate in this scenario will be almost negligible, as evidenced by calculations for CO on Cu in Section S1 of the Supporting Information. Therefore, molecular vibrations will dominantly dissipate energy into EHPs and phonons in the tip. This motivates our model design: We consider four different possible adsorption environments: CO on a sharp tip (sharp tip), CO on a blunt tip (blunt tip), CO on an adatom on a surface slab (slab adatom), and CO on a slab (slab), which are depicted in Figure 1c and Figure S3 in the Supporting Information. The underlying surface, in the case of lateral motion of the tip, only serves as a corrugated energy landscape that triggers molecular vibrations4 and therefore does not have to be explicitly considered in the atomistic model. The “tip” models are therefore represented by two-dimensional periodic surface slab geometries. Further details on the models and the computational methodology are provided in the Methods and Section S1 of the Supporting Information.
@@ 2.2 Vibrational mode analysis
CO adsorbed on a flat metal surface has four distinct types of vibrational modes: frustrated translation (FT), frustrated rotation (FR), surface-adsorbate stretch (SA), and the internal stretch of the CO bond (IS), as illustrated in Figure 2. The FT and FR modes consist of two degenerate modes each.
We determine vibrational energies from second-order force constants (i.e. the Hessian) for all atoms in the molecule and the top layers of the four surface models. Simulations include ten layers of freely movable Cu atoms, spanning hundreds of degrees of freedom. Energies, forces, and second-order force constants are evaluated with a MACE graph neural network model34 trained on first-principles data (see Methods). The MACE potential accurately reproduces the underlying HSE06 data, correctly predicting the adsorption sites: the top site for CO on a (111) surface slab of copper and the hollow site for the adatom in the CO-adatom arrangement (see Section S2 of the Supporting Information). The calculated vibrational frequencies for CO on the Cu(111) slab show good agreement with experimental data, as presented in Table 1. Interestingly, the stiffness of low-energy modes increases as the tip becomes blunter, as indicated in Table 1. Moreover, there is a jump in stiffness when transitioning from CO adsorbed on an adatom to CO adsorbed directly on the surface. HAS experiments estimated the energy of the FT modes on a flat Cu(111) surface to be , while the energy dropped to at step edges of the Cu(211) and Cu(511) surfaces.11 This trend is in good agreement with our finding that more tapered (or corrugated) adsorption environments exhibit lower vibrational energies for the FT mode.
@@
The vibrational density of states (VDOS) (see Figure 3) provides an overview of all vibrational modes. The VDOS is determined from the same molecular dynamics (MD) simulations that yield the results presented in Phonon-phonon coupling. We differentiate between modes localised dominantly at the molecule (orange in Figure 3), where the majority of the kinetic energy is contained in the CO and substrate modes (black in Figure 3), where most of the energy is in the motion of the Cu atoms. The FR, SA, and IS vibrational modes of CO exhibit higher vibrational energy compared to most copper substrate modes, which typically range between approximately and . We note that the VDOS shown here is normalised by the number of atoms. The substrate vibrational states are more localised in energy (or frequency) for slab and slab-adatom surfaces and more delocalised for the sharp and blunt tips. For sharp and blunt tips, a significant portion of the surrounding space is vacuum, whereas this space is occupied by atoms in the slab and slab-adatom surfaces. The presence of more atoms in these flatter adsorption environments results in a greater number of substrate vibrational states per volume. This facilitates stronger overall coupling between the CO and substrate vibrational modes, provided that the CO and Cu degrees of freedom are sufficiently coupled. As a rule of thumb, if the off-diagonal elements of the Hessian (for CO-Cu coupling) are of a similar magnitude, we can expect strong coupling. As shown in Figure S4 in the Supporting Information, this is indeed the case. In the case of the sharp and blunt tip, we find a strong coupling for the CO to a small number of Cu degrees of freedom. The slab adatom and slab geometries exhibit weaker coupling but a larger number of states. When discussing PPC, we will return to this observation.
@@ 2.3 Phonon-phonon coupling
The atomic motion of the CO molecule is coupled to the atomic motion of the Cu substrate through phonon-phonon interactions and to EHP through EPC.2 Even at low temperatures, PPC plays a significant role in the dynamics of molecular vibrations. In our study, we use a temperature of , which is commonly employed in experiments such as AFM measurements with CO-decorated tips. PPC is determined by analysing the dynamics of a CO molecule on adsorption environments where approximately 500 Cu atoms are free to move per unit cell, allowing energy transfer between CO modes and substrate modes. To determine coupling strengths and vibrational lifetimes, we use vibrational analysis of MD trajectories in thermodynamic equilibrium.35 We refer to this approach as equilibrium correlation analysis. A conceptually related approach, which combines vibrational projection, statistical analysis, and local quantum vibrational embedding, was recently introduced to investigate ultrafast energy transfer pathways from lattice phonons to intramolecular modes in molecular crystals.36 We first thermalise each system at a temperature of and then allow it to evolve in thermodynamic equilibrium over time. Over time, the velocity time series of the atoms becomes decorrelated, as the atoms transfer mechanical energy between each other. The rate of this decorrelation can be found in the cross-correlation function. For this analysis, we employ normal mode decomposition to project the velocity time series onto the basis defined by the vibrational eigenvectors. For the velocity time series of two modes and the velocity cross-correlation function is as follows:
@@
The cross-spectrum provides valuable insights into vibrational lifetimes and relative coupling strengths. Let us first discuss the coupling strength, which can be determined by analysing the off-diagonal elements . These exhibit two Lorentzian-shaped peaks at the characteristic frequencies of modes and , with their heights reflecting the coupling strength between these modes (see Section S6.2 of the Supporting Information). To further understand this interaction, we examine the coupling density , defined as the sum of the cross spectra for all modes correlated with the mode of interest :
A broad peak in the coupling density at a given frequency means that modes at this frequency are correlated with the mode of interest, , indicating strong vibrational coupling. Our analysis reveals that the coupling is strongest with adjacent substrate vibrational modes, as shown in Figure 4 and Figures S24, S25, S26, and S27 in the Supporting Information. This effect is less pronounced in case of the energetically decoupled SA and IS modes. The coupling between different CO modes is small in comparison, indicating that intramolecular vibrational redistribution plays only a minor role. Notably, the FR and IS modes exhibit the strongest coupling.
The geometry of the surface has a substantial impact on the coupling density. Among the different geometries, the sharp tip exhibits the weakest coupling, while the slab geometry has the strongest coupling density. This observation can be attributed to the VDOS discussed earlier. As surface geometry becomes more atomically flat, the VDOS increases, offering more vibrational states to accommodate the energy of CO modes and thereby enhancing PPC.